OpenStudy (rachie19):

help with these two probems? plzzz? get the exact and calculator answer 1. log(3)(7x-4)=2 2. 2log(4)3x+log(4)5x=log(4)225

OpenStudy (mathmale):

log(3)(7x-4)=2 can be better written (and be less confusing) if done in Equation (below):\[\log_{3} (7x-4)=2\]

OpenStudy (mathmale):

Since the log and the exponential functions are inverses, \[3^{\log_{3}(7x-4) }\]

OpenStudy (mathmale):

boils down to 7x-4. Now look at the other side of the given equation. It's 2. Use this 2 as the exponent of 3. Note how we're using the same base (3) on both sides of the equation. What do you get?

OpenStudy (rachie19):

I don't understand

OpenStudy (mathmale):

Have you studied the basic properties of exponential and logarithmic functions? Did you know that \[\log_{a} a^x = x?\]

OpenStudy (mathmale):

that's because the log and the exp function are inverses of one another. Before I can help you well, I need to know what you undrstand and what you don't understand about the problem given you here. log(3)(7x-4)=2 should be written as \[\log_{3} (7x-4) = 2. \]

OpenStudy (mathmale):

Use the property of inverse functions to eliminate the operator (command) \[\log_{3} \]

OpenStudy (mathmale):

So again, please tell me as best you can what you do and what you don't understand. We'll go from there.

OpenStudy (unklerhaukus):

1.\[\begin{align} \log_3(7x-4) &= 2\\ 7x-4 &= 3^2\\ 7x &= 3^2+4\\ x &= \frac{3^2+4}7\\ &= \end{align}\]